please help me put the following ellipse equation into standard form.
5x^2+4y^2-30x+8y+29=0

Answer by lwsshak3(11628)

please help me put the following ellipse equation into standard form.
5x^2+4y^2-30x+8y+29=0
complete the square
5(x^2+6x+9)+4(y^2+2y+1)=-29+45+4=20
5(x+3)^2+4(y+1)^2=20
divide by 20
(x+3)^2/4+(y+1)^2/5=1
This is an equation of an ellipse with vertical major axis of the standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1, (a>b), with(h,k) being the (x,y) coordinates of the center